OpenStudy (anonymous):

Find the solution to the differential equation 7(du/dt)=u^2 u(0)=6 U=? Find the solution to the differential equation 7(du/dt)=u^2 u(0)=6 U=? @Mathematics

OpenStudy (anonymous):

so far i have 7/u=t

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

This is a first-order nonlinear differential equation. Let's rewrite it as \[{du \over u^2} = {1 \over 7} dt\]Integrate both sides\[-{1 \over u} = {1 \over 7} t + C_1\] Solve for u\[u = -{7 \over t} + {1 \over C_1} = -{ 7 \over 7C_1 + t}\] Solving for C from initial conditions. \[6 = -{7 \over 7C_1} = -C_1\]

OpenStudy (anonymous):

ok let me do it

OpenStudy (anonymous):

hey man im stuck

OpenStudy (anonymous):

wat was your answer?

OpenStudy (anonymous):

(-7/t)+6?

OpenStudy (anonymous):

\[U = {1 \over 6} - {7 \over t}\]

OpenStudy (anonymous):

You're close.

OpenStudy (anonymous):

my math program says that is wrong

OpenStudy (anonymous):

What answer does your math program give you?

OpenStudy (anonymous):

thats the thing, it does not give you the answer

OpenStudy (anonymous):

let me just post another question and if you find the reason why your other answer could be wrong just message me or somthing

OpenStudy (anonymous):

\[U = {42 \over 7 - 6t}\] Give wolframalpha.com a try.

OpenStudy (anonymous):

ya thats right thank you